The equilibrium constant Kc for the reaction below is 3.75 at 300 K.
What is the value of Kc for the equilibrium below at 300 K?
1/2 A 2 + 1/2 B 2 -------> AB
1
Expert's answer
2019-04-17T04:17:19-0400

Solution.

To find Ks for equilibrium below at 300 K, it is necessary to use the following formula:

"Kc' = Kc\/(RT)^f"

f - the difference between the number of moles of gaseous reaction products and the number of moles of the starting materials.

Kc' = 3.75/(8.31*300)^0 = 3.75

Answer:

Kc' = 3.75/(8.31*300)^0 = 3.75